OpenStudy (anonymous):

7/12 - 1/4 and 5/6 - 3/8 and 23/24 - 7/8 in fractions

hero (hero):

one question at a time mate

OpenStudy (anonymous):

k bro

OpenStudy (anonymous):

7/12 - 1/4 = 1/3 5/6 - 3/8 = 11/24 23/24 - 7/8 = 1/12 Do you not know how to do these or something? You keep asking for help with subtracting fractions and it's not really that hard. You should start practicing on your own!

hero (hero):

Some people have trouble with math because their minds cannot think computationally

OpenStudy (anonymous):

me neither man

hero (hero):

I didn't say me silly.

OpenStudy (anonymous):

also sometimes i just ask to confirm if its right

OpenStudy (anonymous):

Ah, I see. Sorry if I offended you or anything, just wanted to make sure that you actually know how to do it so you won't end up stressing when you find something like this on your test :)

OpenStudy (anonymous):

its ok you didnt offend me

hero (hero):

Zacho, if you're going to post something like that, you should include all your steps

hero (hero):

Posting the question, then the answer doesn't really help much

OpenStudy (anonymous):

what do you mean

OpenStudy (anonymous):

oh i get it

hero (hero):

Ever heard of intermediary steps?